OpenStudy (anonymous):

medals rewarded! A Super Bounce Ball is dropped from a height of 64 ft. With each bounce, the ball reaches a height that is three-fourths the height of the previous bounce. After how many bounces will the ball bounce up to a height less than 9 in.? a.3 b.5 c.4 d.6

OpenStudy (anonymous):

one bound \(64\times \frac{3}{4}\)

OpenStudy (anonymous):

two bounces \(64\times (\frac{3}{4})^2\)

OpenStudy (anonymous):

etc